OpenStudy (anonymous):

How do I solve cube root of 2 to the sixth power, to the x power

OpenStudy (anonymous):

i cant understand it does it really say \[\left(\sqrt[3]{2^6}\right)^x\]?

OpenStudy (anonymous):

one thing is for sure \[\sqrt[3]{2^6}=2^2=4\] after that i guess you have \(4^x\)

OpenStudy (anonymous):

Sorry I'm not good with computers. But yeah that's what I meant